OpenStudy (anonymous):

the area of a rectangle is given by the equation w(2w+1)=36 what is the width of the triangle?

OpenStudy (anonymous):

First factor out w(2w+1) to get 2x^2+w=36

OpenStudy (anonymous):

ok I know the answer is 4 but I need to work it out

OpenStudy (anonymous):

\[2w^2+w=36\] Take square root: \[\sqrt{36}=6\] \[2w+w=6\]

OpenStudy (anonymous):

\[3w=6\] \[w=2\]

OpenStudy (anonymous):

but the answer should be 4 so how is it 2

OpenStudy (anonymous):

I showed you how.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Does it have answer choices and your welcome.